OpenStudy (pphalke):

If a right triangle has legs measuring 6 feet and 8 feet, how long is the hypotenuse?

Vocaloid (vocaloid):

leg^2 + leg^2 = hypotenuse^2

Vocaloid (vocaloid):

6^2 + 8^2 = c^2 c = ?

OpenStudy (pphalke):

100